### v1

# Clothing agnostic Pre-inpainting Virtual Try-ON

Abstract
With the development of deep learning technology, virtual try-on technology has become an important application value in the fields of e-commerce, fashion, and entertainment. The recently proposed Leffa has improved the texture distortion problem of diffu-sion-based models, but there are limitations in that the bottom detection inaccuracy and the existing clothing silhouette remain in the synthesis results. To solve this problem, this study proposes CaP-VTON (Clothing agnostic Pre-inpainting Virtual Try-ON). CaP-VTON has improved the naturalness and consistency of whole-body clothing syn-thesis by integrating multi-category masking based on Dress Code and skin inpainting based on Stable Diffusion. In particular, a generate skin module was introduced to solve the skin restoration problem that occurs when long-sleeved images are converted into short-sleeved or sleeveless ones, and high-quality restoration was implemented consider-ing the human body posture and color. As a result, CaP-VTON recorded 92.5%, which is 15.4% better than Leffa in short-sleeved synthesis accuracy, and showed the performance of consistently reproducing the style and shape of reference clothing in visual evaluation. These structures maintain model-agnostic properties and are applicable to various diffu-sion-based virtual inspection systems, and can contribute to applications that require high-precision virtual wearing, such as e-commerce, custom styling, and avatar creation.
1 Introduction
With the rapid development of computer vision and deep learning technology, Con-trollable person image generation models have come to have important application value in the fields of e-commerce, fashion industry, and entertainment. In particular, virtual try-on technology is contributing to improved purchase satisfaction and reduced return rates by allowing consumers to check the wearing effect in advance without actually wearing clothes in an online shopping environment. Against this background, the Learn-ing Flow Fields in Attention(Leffa) method proposed by Zhou et al.[1] is drawing attention as an innovative approach to solving the detailed texture distortion problem that existing diffusion-based models were experiencing.
Existing controllable person image generation methods have achieved a high level of overall image quality, but have still shown limitations in accurately preserving fine tex-ture details from reference images. These issues have been particularly noticeable in clothing containing complex patterns, logos, and texts, and have acted as important bar-riers to implementing practical virtual inspection services. Leffa identified the root cause of this problem as the attention mechanism’s inability to focus on the correct area of the reference image, and addressed this by introducing a normalization loss that induces the learning of the flow field within the attention layer to focus explicitly on the correct refer-ence area.
Figure 1: Limitations of previous Leffa model: (a) Problem of detecting bottom, (b) Synthesis limitations dependent on source input images
However, there were synthesis results that could not be covered by the existing Leffa model, and we supplemented them by adding several things from the pipelines of the ex-isting study. The points that can be supplemented by the existing Leffa model are shown in Figure 1. The first can be confirmed in Figure 1(a). Since the existing Leffa model is mainly optimized for virtual inspection centered on upper-body clothes, detection and masking of the lower-body areas are relatively vulnerable. This can put great restrictions on this part in a virtual inspection scenario centered on whole-body clothing or bottoms. In particular, when the lower-body could not be detected in the input image, there was a limitation in that the whole body was synthesized even if the upper-body clothes were synthesized. Therefore, a separate specialized preprocessing strategy that elaborately re-flects this is needed. The second was that the existing clothing characteristics included in the input source image had more than a certain level of influence on the final output result. Figure 1(b) was outputted as a Leffa-Corrective Masking version that improved the first limitation to confirm independently of the first problem. When trying to synthesize short-sleeved clothing with the image of a person wearing a long-sleeved top as shown in Figure 1(b), it was observed that the existing long-sleeved shape was maintained regard-less of the input garment. This seemed to be because the skin area to be exposed, such as the forearm, shoulder, and neck, was not sufficiently restored in the process of removing clothing, and the existing clothing type acted as an interference factor in the subsequent clothing synthesis process. This phenomenon could lead to unnatural results in the syn-thesis between clothing with large differences in shape such as sleeve length and neckline.
By integrating the Dress Code-based improved masking mechanism and skin inpainting pipeline while maintaining the advantages of the existing Leffa pipeline, this study aims to achieve complete synthesis results in all clothing categories and bottoms. This, combined with Leffa’s excellent attention-based flow field learning ability, was able to present an integrated solution that satisfies the three goals at the same time: detailed texture preservation, full-body clothing synthesis, and input image independent clothing replacement. In particular, it was confirmed that the interference phenomenon caused by existing clothing characteristics was significantly improved along with the morphological inconsistency, the unnaturalness of the boundary, and the incongruity with the body ratio shown in the bottom synthesis. In addition, this study is of high practical value in that it provides a universal improvement method that can be applied to various diffusion-based models through improved accuracy of the masking process and improved skin restoration quality while maintaining the model-agnostic characteristics of the existing Leffa model. This approach is differentiated from existing studies that have presented solutions de-pendent on specific models or datasets, and is expected to be widely used in various ap-plications in the virtual try-on field in the future.
2 Related Work
2.1 Evaluation of Generative Models
The generative model is difficult to evaluate compared to the classification model. This is because there is no single image to compare with the results obtained through the generative model. It is necessary to create an index that can evaluate following the distri-bution of actual domain images, and various evaluation indicators have been proposed. The indicators used in the evaluation are as follows.
FID(Frechet Inception Distance) is a key indicator of the quality of images created by generative models[2]. FID measures the statistical distance between the two distributions by comparing the distribution of real and generated images, extracts feature vectors in 2,048 dimensions using a pre-trained Inception v3 model, and measures the distance by calculating the mean ($\mu$) and covariance (C) of the real and generated images. This means that the lower the FID value, the more similar the generated image is to the real image.
$d^{2}=\|\mu_{1}-\mu_{2}\|^{2}+\operatorname{Tr}\left(C_{1}+C_{2}-2\left(C_{1}C_{2}\right)^{1/2}\right)$
SSIM(Structural Similarity Index Measure) is an index that quantifies structural sim-ilarity between two images[3]. SSIM outputs a value between 0 and 1 by comparing the luminance, contrast, and structural elements of two images, and a closer to 1, the more structurally similar the two images are. In particular, SSIM is more advantageous in eval-uating the degree of retention of patterns and local structures than errors in pixel units. Generally, it is used to evaluate the visual quality of high-resolution images.
$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$
LPIPS(Learned Perceptual Image Patch Similarity) is an indicator that quantifies perceptual differences between two images using deep learning-based image recognition models[4]. LPIPS measures the similarity by extracting feature vectors from different lay-ers of pre-trained neural networks (e.g., AlexNet, VGG, SqueezeNet, etc.). The lower the LPIPS value, the more visually similar the two images are, and it has a complementary relationship with FID and SSIM in that it better reflects high-dimensional semantic simi-larity than simple pixel differences.
$\mathrm{LPIPS}(x,y)=\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}\left\|w_{l}\odot\left(\hat{f}_{l}^{x}(h,w)-\hat{f}_{l}^{y}(h,w)\right)\right\|_{2}^{2}$
The various indicators described in this chapter are used to compare the results of the model we proposed in 4.2.2 with existing models. In addition to this, in this study, the form-consistency of the virtual inspection model is regarded as a key criterion directly re-lated to the actual service quality, and a statistical-based evaluation index was introduced to measure this objectively. The normal output rate is a representative statistical-based quantitative indicator, which refers to the percentage (%) of cases where the normal short-sleeved silhouette is accurately synthesized, especially when the image of wearing long-sleeve clothing is input and the reference clothing is short-sleeve. The main purpose of this indicator is to directly quantify the degree of reflection of the reference clothing sil-houette and the interference phenomenon of input clothing characteristics, which are dif-ficult to explain with simple image quality and similarity indicators (FID, SSIM, LPIPS, etc.). Existing models often did not match the actual reference clothing type (e.g., short-sleeve, sleeveless, etc.) and left the input clothing silhouette (especially long-sleeve), but the proposed technique was fundamentally improved through pipeline strategies such as generate skin and improvement masking. The normal output rate calculation is performed by evaluating the output image when the reference clothing is short-sleeved for long-sleeved input images in the dataset, classifying whether the short-sleeved form is actually implemented (normal/abnormal) as in the example, and then calculating the normal ratio.
2.2 Virtual Try-on Models
Virtual Try-on models can be divided into classical models, diffusion-based models, and attention-based models. Early Virtual Try-On approaches mainly used GAN-based warping techniques and segmentation. CP-VTON[5] handled clothing deformation using thin plate spline (TPS) warping and segmentation, but there was a limit to realistic image generation as the FID value was very high at 47.36. Since then, VITON-HD[6] has achieved FID 11.74 at 1024×768 resolution by introducing alignment-Aware segment (ALIAS) nor-malization and warping unit, but there was still a limit to accurate clothing alignment and natural deformation in complex poses. ACGPN[7] introduced an adaptive content generating and preserving network to record an FID of 26.45, but its limited performance at 256×192 resolution made it difficult to expand high resolution, and HR-VITON[8] im-proved to FID of 10.91 afterwards, but the sleeve-squeezing and waist-squeezing artifact problems persisted. These classical models show an average FID of 24.06, with funda-mental limitations in preserving complex clothing textures and generating realistic wear-ing effects.
With the advent of diffusion models from 2023, great progress has been made in the field of Virtual Try-On. Diffusion models enabled more stable learning and high-quality image generation compared to GAN(Generative Adversarial Network)[9], and achieved a lot of performance improvement compared to the past with an average FID of 7.18. Sta-bleVITON[10] achieved FID 6.52 by applying zero cross-attention to LDM (Latent Diffu-sion Model), but there was a segment detail loss problem due to attrition misalignment. IDM-VTON[11] improved to FID 6.29 by utilizing segment-agnostic person representation and mask, but there was a limit to customization deterrence and continuous consistency in complex clothing transformation. LaDI-VTON[12] recorded a relatively high fid value with an FID of 8.85 by combining Latent Diffusion, Textual-Inversion, and CLIP. OOTDif-fusion [13] implemented multiple outfit processing through parallel outfitting UNet and outfitting fusion, and recorded an average fid score of about 10 with 11.03 on top, 9.72 on bottom, and 10.65 on dress for each part.
Later, the introduction of transformers and attention mechanisms enabled more so-phisticated feature matching and spatial relationship modeling even in Virtual Try-On. The multi-head self-attention proposed in Vaswani, A. et al.[14] became the basis of the transformer model, and models that use it in Virtual Try-On also began to emerge. Cat-VTON[15] selectively fine-tuned only the self-attention module within the diffusion U-Net and conducted lightweight learning with 49.57M parameters, removing text encoders and cross-attention, and maximizing efficiency by simply concatenating and inputting human and costume images in the spatial dimension. However, detailed semantic alignment control was limited. Leffa[1] improved the self-attention and cross-attention map by ap-plying flow field regularization to the attachment of the diffusion model. This enabled more sophisticated attention control than the existing mask inpainting method, zero cross-attention method, and prompt-aware mask method, but there were still problems with the complexity and fine-detail distortion of flow field regularization.
Existing Leffa applied flow field regularization unique to the attention mechanism to enable more sophisticated attention control than existing methods, but there are still limi-tations in complex clothing deformation, detailed texture preservation, and accurate area segmentation in the pre-processing stage. This was a situation as shown in Figure 1, and several steps were added to the model’s pipeline to overcome this limitation.
In this study, we propose an inpainted pre-processing input-based attention flow system to improve the limitations found by existing Leffa model. We propose the CaP-VTON(Clothing agnostic Pre-inpainting Virtual Try-ON) model by adding a part that can compensate for the shortcomings to the existing VITON-HD model of the Leffa model. Through this model, the stability of the diffusion model, the attention mechanism of the transformer, and the advantages of the classical warping technique are selectively integrated to solve the problems of the limitations of the attention control, the inaccuracy of the pre-processing step, and the lack of consistency for various clothing types of existing studies.
Figure 2: Proposed method: CaP-VTON(Clothing agnostic Pre-inpainting Virtual Try-ON)
3 Proposed Method
3.1 System Flow: Concept of CaP-VTON
This study proposes the CaP-VTON(Clothing agnostic Pre-inpainting Virtual Try-ON) model, a new model that integrates two key improvement factors to complement the structural limitations of the existing Leffa pipeline and enable a natural and coherent transition between various clothing categories. The proposed improvement factors are the Dress Code-based multi-category masking stage, which covers full-body clothing such as top, bottom, and dress, and the Stable Diffusion-based skin inpainting pre-processing stage for removing existing clothing silhouettes and restoring exposed skin.
The overall flow is shown in Figure 2. This study can be largely divided into two stages. One is the Modified Leffa Pipeline for the limitation that detection and masking of the lower-body cloths area is relatively weak because the existing VITON-HD Leffa model is mainly optimized for virtual inspection centered on upper-body cloths, and the other is the Stable Diffusion Generation Skin Inpainting step for the limitation that the existing clothing characteristics included in the input source image affect the final output result more than a certain level. In the first step, when the user inputs a source image, the se-mantic part of the human body is first subdivided using the SCHP(Self-Correction for Human Parsing) model[16], and at the same time, various clothing areas such as top, bot-tom, and dress are accurately masked through the Dress Code model. In addition, by ex-tracting the person’s posture information using OpenPose[17], the human structure is not distorted during the subsequent inpainting process. The body mask and pose information obtained in this step are used in the inpainting process to remove the existing clothing ar-ea and naturally restore the skin and short-sleeved silhouette. In the first step, an pre-processed inpainting image(B in Figure 2) is generated from the input source image(A in Figure 2). In the second step, a mask of the pre-processed image is made through the Dress Code model by using the pre-processed inpainting image(B in Figure 2) output in a clean form without the silhouette of the existing long-sleeved clothing, and is transmitted to diffusion painting applying the attachment-based flow field regularization of the VI-TON-HD model along with the reference clothing information to generate a final virtual fitting image(C in Figure 2).
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 3: Changes in output results according to Leffa clothing type.
3.2 Stable Diffusion Generate Skin Inpainting: Removing existing clothing silhouette
When trying to synthesize short-sleeved clothing with the image of a person wearing a long-sleeved top in the existing model, there was a problem that the existing long-sleeved shape was maintained regardless of the type of reference clothing as shown in Figures 3(A) and (C), and the problem that the silhouette of the existing clothing re-mained as shown in Figure 3(B). This was because the silhouette of the existing clothing interfered with the synthesis of new clothing because the exposed skin areas such as the forearms, shoulders, and neck were not properly restored during the clothing removal process.
In order to essentially solve this problem, the generate skin method pipeline has adopted an approach that inserts short-sleeved clothes and realistic skin into the area where existing clothes have been removed, which does not affect the output. This was in-tended to accurately reflect the unique shape and style of the reference clothes regardless of the existing clothing characteristics of the input image.
3.2.1 Generating input information to build synthetic model
In order to perform inpainting through the Generate Skin function, an OpenPose im-age containing the inpainting mask image and the pose information of the input image is required. The Generate Skin Inpainting Mask Output of Figure 4 is generated using the masking pipeline provided by Leffa’s DressCode model. In addition, the OpenPose Output of Figure 5 is used as conditioning information of the ControlNet in the process of Stable Diffusion-based inpainting. As shown in Figure 5, when the ControlNet is not used, the pose identity of the original person is not accurately transmitted to the network, resulting in a problem of creating a structurally distorted image, such as a misalignment of the body position such as the arm or shoulder or mixing with the background. As a result, we determined that the problem of preserving the pose and securing the accuracy of the skin area in the process of inpainting directly affects the realism of the final result, and to solve this problem, the Generate Skin Inpainting Mask and OpenPose-based ControlNet[18] in-formation were used as input information for the Generate Skin function. In addition, as can be seen in Figure 5, there were cases where the color of the inpainted skin did not properly reflect the skin tone of the original image, resulting in the problem of creating darker or brighter skin than it actually is. To compensate for this, the Skin Tone Extraction technique was introduced to create a skin texture that naturally harmonizes with the original image. This technique is implemented by detecting the skin area in the YCrCb color space and calculating the average skin tone value in the HSV color space. The skin area created through this is naturally blended with the skin tone of the original image, so that the overall visual consistency of the resulting image can be maintained.
Figure 4: Preprocessed Images for Stable Diffusion Generate Skin inpainting
Figure 5: Comparison of results with/without Openpose ControlNet
3.2.2 MajicMix Realistic model based high quality skin synthesis
The core of the Generate Skin pipeline is inpainting the skin area using the MajicMix Realistic model[19]. This model is based on Stable Diffusion 1.5, and is a high-quality checkpoint model trained by fusing various model weights. In particular, it is optimized for generating Asian skin tones, so it has a strength in deriving natural results from Asian images. In addition, this model introduces Noise Offset Technique, which enables more sophisticated light source expression and smooth shadowing during the diffusion process. It exhibits excellent performance in restoring texture, shade, and color of complex skin ar-eas such as arms, shoulders, and neck exposed after clothing removal. Figure 6 shows the results output through the generate skin pipeline.
Figure 6: Examples of results of generate skin pipeline
3.3 Modified Leffa pipeline: Multi-category enabled based on DressCode
There are two models in the Leffa model: one trained with VITON-HD data and one trained with DressCode data. The VITON-HD Leffa model was optimized for phase-oriented synthesis. For this reason, there were limitations in the detection of the lower area and the whole body synthesis, and the image with a high proportion of the lower body showed unstable synthesis results. To solve this problem, as shown in Figure 7, the inpainting Mask used the DressCode model, and the one of the VITON-HD model with high output image quality was ensembled in the Leffa Diffusion inpainting output. A masking strategy was constructed using the multi-category structure of the DressCode dataset.
Figure 7: Modified Ensemble model of VITON-HD and DressCode Model
3.3.1 Characteristic of VITON-HD and DressCode Model
Figure 8 shows the comparison of the masking quality of the two models. The VI-TON-HD model is trained on the VITON-HD dataset specializing in upper body clothing images and is mainly optimized for image synthesis. Accordingly, the image centered on the upper body shows excellent visual quality, but the information on the lower body is not included, so there is a limit to the detection and masking of the lower body area. On the other hand, the Dress Code model was learned based on a dataset including various categories of clothing such as tops, bottoms, and dresses, which allows masking for full-body clothing. In fact, in this study, by using the Dress Code model when generating an inpainting mask, stable masking results were secured in multi-category situations.
Figure 8: Comparison of two masking methods: Red colored highlighted regions for masking
To compensate for these structural differences, this study designed a pipeline in a way that combines the advantages of the two models. In the in-painting masking process, multi-category detection was performed using the Dress Code model, and better image quality was secured by using the VITON-HD model in the final Diffusion synthesis step. To verify this numerically, the synthesis performance of the two models was compared as shown in Table 1. As a result, the FID score was 4.54 for the VITON-HD model, which was lower than the Dress Code model (5.05), showing an advantage in terms of the overall visual quality of the generated image. On the other hand, in the SSIM and LPIPS indicators, the Dress Code model is better, showing that it is superior in terms of structural and de-tailed texture expression of actual clothes.
Table 1: Numerical Comparison between two models.
Model FID $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
VITON-HD 4.54 0.899 0.048 8.52
DressCode 5.05 0.949 0.021 10.73
4 Experimental Results
4.1 Experimental Environment
This study was experimented in the Google Colab environment, and Ubuntu 20.04 LTS was used as an operating system. The programming language consisted of Python 3.11.13 version, and the compiler was GCC 11.4.0. As the main library, packages related to image processing and deep learning such as OpenCV and TorchGeometry were used, centering on PyTorch (version 1.6.0 and higher, GPU acceleration support). Stable Diffu-sion was used as a model for image generation, and the image was created in a 30-step setting based on the Diffuser library or similar framework. The experimental environment was operated through the Colab interface based on Jupiter Notebook. The hardware con-figuration consisted of NVIDIA A100 GPU (40GB memory), 83.5GB of system RAM, and Intel Xeon-based CPU, and the disk space used in the experiment was about 80GB. When Stable Diffusion was configured to a 30-step setting, it took about 8 seconds on average to generate a single image in the NVIDIA A100 GPU environment. The data using the model production and evaluation were made using the Viton-HD dataset, 1100 clothing images, and 1100 people images. The person used to create the resultant image was produced through a generative model.
4.2 Results
4.2.1 Visual Comparison of the synthesis results
Figure 9 is a visual comparison of the synthesis results generated by various virtual wearing models based on the same input image and the same reference clothing. This comparison was evaluated beyond simple visual quality, focusing on whether the silhou-ette of the reference clothing (especially the sleeve length and arm exposure area) was ac-curately reflected.
As a result of the comparison, the existing CATVTON, IDM-VTON, Leffa-VITONHD, and Leffa-DressCode models were commonly influenced by the existing long-sleeved sil-houette of the input image, showing limitations in not being able to accurately reproduce the design of reference clothing, such as abnormally long sleeves in the output image or insufficient exposure of arms and shoulders. In particular, it was observed that the Leffa-VITONHD model was synthesized up to the lower half, which should not be syn-thesized, and despite the fact that short-sleeved clothing was actually applied, it was ob-served that the shape of the existing long-sleeved sleeves remained on the output or the human body was unnaturally distorted.
Figure 9: Result of our method (on last column): 1st column is input person image, and 2nd images are the clothes for change. 3rd–7th columns are synthesized results of other methods.
On the other hand, the proposed CaP-VTON (Ours) model effectively removes exist-ing long-sleeved clothing and naturally and consistently restores skin areas such as arms, shoulders, and armpits to be exposed by utilizing the generate skin-based preprocessing technique. As a result, the short-sleeved silhouette of the reference clothing was clearly implemented in the output image, and the pattern and shape were neatly reflected with-out colliding with the existing clothing. This difference does not stop at just the visual naturalness, but acts as an important indicator for evaluating how accurately the style of the reference clothing can be reflected. While existing models are strongly influenced by input clothing, CaP-VTON shows that it is possible to reproduce a form independent of input clothing through structural consistency-based preprocessing. This is a critically important factor in the whole-body virtual wearing system for which practical clothing replacement is the purpose, and this study is one of the key contribution points that dis-tinguishes it from the existing methodology.
In conclusion, CaP-VTON is a model that most accurately reflects the style and sil-houette of reference clothing, experimentally demonstrating that it provides the most reli-able synthetic quality from a user experience perspective beyond simple image quality or structural similarity indicators.
4.2.2 Quantitative Comparison
Table 2 is the result of comparing the performance between the existing virtual wear-ing models and the proposed CaP-VTON model with FID, SSIM, and LPIPS indicators based on the category of the award. The proposed CaP-VTON model showed a slightly lower quantitative performance with FID 11.46, SSIM 0.8573, and LPIPS 0.0849 due to the additional image inference pipeline(Generate Skin Method).
However, these figures are general indicators focusing on overall image quality and perception-based similarity, and there is a limit to evaluating the accuracy of conversion from long-sleeve to short-sleeve, which is the purpose of this study. Therefore, observa-tion-based accuracy was added and evaluated.
Table 2: Numerical comparison with other models. The lower the FID and LPIPS, the better the quality, and the higher the SSIM, the better the quality.
Model paired unpaired
FID $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
CatVTON 5.42 0.870 0.057 9.02
IDM-VTON 5.76 0.850 0.063 9.84
OOTDiffusion 9.30 0.819 0.088 12.41
Leffa 4.54 0.899 0.048 8.52
CaP-VTON (ours) 6.55 0.897 0.062 11.28
Figure 10: Comparison of short sleeve & sleeveless clothing output
Figure 11: Comparison the percentage of correct ouput
4.2.3 Evaluation based on Short-Sleeve Synthesis Probability
This study additionally conducted a quantitative experiment to evaluate whether or not the reference clothing was actually synthesized as a normal short sleeve when it was short-sleeved, targeting the long-sleeved wearing input image. This is to quantitatively verify the form-consistency of the model’s clothing shape by checking how much the sil-houette of the existing clothing interferes with the final output. As shown in Figure 10, it can be seen that the Leffa model is very vulnerable to this part. Figure 11 is a visualization of the results of classifying the output results into normal output (when a short-sleeved silhouette matching the reference clothing is accurately generated) and abnormal output (when a long-sleeved silhouette remains or short-sleeved clothing is synthesized as long-sleeved clothing).
This result means that CaP-VTON recorded about 15.4%p higher accuracy of clothing silhouette consistency than Leffa. In particular, CaP-VTON showed excellent performance in the ability to accurately reflect short-sleeved reference clothing after removing long sleeves and naturally restoring the skin of the arms and shoulders, which can be said to be the result of empirically proving the effectiveness of generate skin-based pre-processing and improved masking pipeline.
On the other hand, Leffa showed excellent quantitative performance in terms of FID and SSIM, but the above experimental results show the limitation that it is difficult to fully explain the suitability of the model’s clothing replacement only with quantitative image quality evaluation. In actual applications, the consistency-based evaluation method pro-posed in this study is a very important measure in a practical sense because the accuracy and visual naturalness of the clothing shape conversion desired by the user act as more important factors.
In conclusion, CaP-VTON showed a distinct improvement over Leffa in terms of structural consistency, meaning that the pipeline was designed to minimize dependence on existing clothing types and accurately reflect the style of reference clothing. These re-sults support the feasibility of CaP-VTON in high-dimensional applications such as full-body virtual wear and style transformation.
5 Discussion
5.1 Improved entire pipeline
The improved Masking Pipeline and Generate Skin method-based skin inpainting pipelines do not just complement each function, but also create complementary synergies at the entire pipeline level. In particular, the organic integration of these two modules ef-fectively overcomes the structural limitations of the existing Leffa model, and as a result, natural and high-quality results can be stably produced even in a full-body clothing re-placement scenario.
Since the existing Leffa model specialized in top-centered synthesis, it showed a lim-itation in not being able to properly handle complex clothing structures such as bottoms or dress categories. These limitations result in two key problems: incompleteness of detec-tion of clothing areas and the remaining of existing clothing silhouettes. Therefore, this study introduced a double complementary pipeline structure that considers both the im-provement of masking accuracy and the effect of removing existing clothing.
Masking Pipeline utilizes multi-category information from the Dress Code dataset to perform accurate clothing boundary detection, and based on the results, generates an ap-propriate inpainting mask to remove existing clothing areas. Subsequently, the generate skin pipeline performs a function of blocking existing clothing characteristics included in the input image from interfering with the final output by realistically restoring the skin and short-sleeved silhouette to the removed area. This dual processing structure provides an ideal condition for accurately and consistently reflecting the shape of the reference clothing regardless of the shape of the existing clothing.
For example, in a scenario where a person wearing a long-sleeve shirt wears a sleeveless dress, the Masking Pipeline accurately detects the long-sleeve area from the shoulder to the wrist, and the Skin Inpainting Pipeline naturally restores the skin in the area, thereby accurately reproducing the exposure design of the reference dress. This is not just a partial modification, but a structural improvement of the entire process leading to pre-masking-inpainting-diffusion for whole-body synthesis.
5.2 Supporting for replacement across different clothing types
This integrated pipeline is designed based on the multi-category structure of the Dress Code dataset, enabling free clothing replacement between two categories: top/bottom, top/dress, and bottom/dress. This function was technically impossible in the top-centered synthesis structure of the existing Leffa model, and it is difficult to implement without the structural improvement proposed in this study. In addition, for complex structural clothing in which the top and bottom are integrated, such as dresses, bottom masking and upper body skin restoration are applied in parallel, enabling natural and flawless synthesis results. This allows users to attempt complex styles that include not only style changes within a single category but also shape transitions between categories, which greatly expands the scope of use of virtual wearing systems.
5.3 Maintaining Model-agnostic properties
On the other hand, another important contribution of this study is that this structural improvement is not dependent on a specific model and maintains the possibility of mod-el-agnostic expansion. Since the proposed improvement methods are not deeply coupled to the internal structure of the Leffa model, but are designed with a modular structure ap-plied to the input image preprocessing and masking generation process, they can be easily integrated into similar diffusion-based models (e.g., HumanDiff, FashionStyleGAN, VI-TONDiffusion, etc.).
In other words, improved masking strategies and generate skin inpainting tech-niques can serve as universal complementary techniques that can dramatically improve the accuracy and style reflection capabilities of existing synthetic models, providing high practical value across virtual wearing techniques.
6 Conclusion
This study proposed a new pipeline structure, CaP-VTON (Clothing agnostic Pre-inpainting Virtual Try-ON), to overcome the limitations of the existing Leffa-based virtual try-on system. While Leffa showed excellent performance in detailed texture repre-sentation through attention-based flow field learning, it had structural limitations such as bottom detection inaccuracy and existing clothing silhouette remaining. In order to solve these two key problems, this study designed a dual complementary pipeline that inte-grates (1) a multi-category masking pipeline based on Dress Code and (2) a generator skin preprocessing mechanism based on Stable Diffusion.
The masking pipeline has been improved to accurately detect various clothing cate-gories such as top, bottom, and dress, and the skin inpainting pipeline is configured to remove existing clothing areas and naturally restore skin and short-sleeved silhouettes to accurately reflect the style of reference clothing. In particular, the generate skin method contributed to the stable transformation from long-sleeved to short-sleeved or sleeveless clothing by naturally restoring skin areas such as arms, shoulders, and armpits according to the human body’s posture after removing clothing. As a result of the experiment, CAP-VTON had the same similar quantitative index compared to the existing Leffa, and at the same time, achieved an accuracy of 92.5% in the evaluation of consistency based on the accuracy of short-sleeved silhouette synthesis, achieving 15.4%p higher performance than Leffa’s 77.1%. In addition, it was confirmed that CAP-VTON consistently reproduces the sleeve length, pattern, and style of reference clothing, unlike other existing models.
This study has the following contributions to the virtual inspection model. First, the bottom and dress masking functions were successfully implemented to support whole-body synthesis. Second, an input-independent synthesis structure was established by removing the existing clothing interference. Finally, the model-agnostic modular structure secured versatility applicable to various diffusion-based models. This contribu-tion can be used as a foundation technology to build a more precise and practical virtual start-up system in the actual e-commerce environment in the future and has the potential to expand into a variety of fields such as simulation of customized clothes, avatar genera-tion, and clothing design preview. Future research is expected to further expand the prac-ticality and applicability of the CAP-VTON pipeline by further elaborating the processing of complex poses, backgrounds, interactions with non-clothing elements such as accesso-ries and hair, improving inference speed in real-time environments, and integrating it with a dynamic clothing recommendation system based on user feedback.